Hart Enterprises recently paid a dividend, D0, of $3.25. It expects to have nonconstant growth of 20% for 2 years followed by a constant rate of 5% thereafter. The firm's required return is 19%. What is the intrinsic value today?
Answer #1

D1=(3.25*1.2)=3.9

D2=(3.9*1.2)=4.68

Value after year 2=(D2*Growth rate)/(Required rate-Growth rate)

=(4.68*1.05)/(0.19-0.05)

=35.1

Hence  intrinsic value today=Future dividend and value*Present value of discounting factor(rate%,time period)

=3.9/1.19+4.68/1.19^2+35.1/1.19^2

=$31.37(Approx).